# thanks: Corresponding authorthanks: Corresponding author

### v1

@@ Abstract
One of the most fundamental relationships in modern cosmology is the cosmic distance duality relation (CDDR), which describes the relationship between the angular diameter distance ($D_{\rm A}$) and the luminosity distance ($D_{\rm L}$), and is expressed as: $\eta(z)=D_{\rm L}(z)(1+z)^{-2}/D_{\rm A}(z)=1$. In this work, we conduct a comprehensive test of the CDDR by combining baryon acoustic oscillation (BAO) data from the SDSS and DESI surveys with type Ia supernova (SN) data from PantheonPlus and DESY5. We utilize an artificial neural network approach to match the SN and BAO data at the same redshift. To explore potential violations of the CDDR, we consider three different parameterizations: (i) $\eta(z)=1+\eta_{0}z$; (ii) $\eta(z)=1+\eta_{0}z/(1+z)$; (iii) $\eta(z)=1+\eta_{0}\ln(1+z)$. Our results indicate that the calibration of the SN absolute magnitude $M_{\rm B}$ plays a crucial role in testing potential deviations from the CDDR, as there exists a significant negative correlation between $\eta_{0}$ and $M_{\rm B}$. For PantheonPlus analysis, when $M_{\rm B}$ is treated as a free parameter, no evidence of CDDR violation is found. In contrast, fixing $M_{\rm B}$ to the $M_{\rm B}^{\rm D20}$ prior with $-19.230\pm 0.040$ mag leads to a deviation at approximately the $2\sigma$ level, while fixing $M_{\rm B}$ to the $M_{\rm B}^{\rm B23}$ prior with $-19.396\pm 0.016$ mag remains in agreement with the CDDR. Furthermore, overall analyses based on the SDSS+DESY5 and DESI+DESY5 data consistently show no evidence of the deviation from the CDDR.
@@ I Introduction
The cosmic distance duality relation (CDDR), also known as the Etherington relation Etherington (1933), is a fundamental relationship in modern cosmology. This relationship is predicated on three fundamental assumptions: the description of space-time by the metric theory of gravity, the photons propagate along null geodesics, and the conservation of the photon number Bassett and Kunz (2004). Specifically, for a given redshift, the relation can be expressed as $D_{\rm L}(z)=(1+z)^{2}D_{\rm A}(z)$, which establishes a connection between two critical measures of cosmic distance for an astronomical object: the luminosity distance ($D_{\rm L}$) and the angular diameter distance ($D_{\rm A}$). Therefore, testing the validity of the CDDR is essential for verifying the theories of gravity and the foundations of modern cosmology Avgoustidis et al. (2010); Gonçalves et al. (2012).
@@
In recent years, a considerable amount of research has employed various astronomical observations to test the CDDR Uzan et al. (2004); De Bernardis et al. (2006); Li et al. (2011); Nair et al. (2011); Holanda et al. (2010); Liang et al. (2013); Meng et al. (2012); Khedekar and Chakraborti (2011); Goncalves et al. (2012); Lima et al. (2011); Cardone et al. (2012); Yang et al. (2013); Wu et al. (2015); Ma and Corasaniti (2018); Holanda et al. (2017); More et al. (2016); Rana et al. (2017); Li and Lin (2018); Lin et al. (2018); Qi et al. (2019); Holanda et al. (2019); Zhou et al. (2021); Qin et al. (2021); Bora and Desai (2021); Mukherjee and Mukherjee (2021); Liu et al. (2021); Xu et al. (2022); Tonghua et al. (2023); Yang et al. (2024); Tang et al. (2025); Jesus et al. (2025); Qi et al. (2025); Alfano and Luongo (2025). Type Ia supernova (SN) have been widely used to determine the luminosity distance, whereas the angular diameter distance, is typically obtained through multiple observational techniques, including (but not limited to) the Sunyaev–Zeldovich effect and gas mass fraction measurements in galaxy clusters Holanda et al. (2010); Li et al. (2011), strong gravitational lensing systems Liao et al. (2022); Ruan et al. (2018); Qi et al. (2025), the angular size of ultra-compact radio sources Li and Lin (2018), and baryon acoustic oscillations (BAO) Wu et al. (2015). In particular, SN and BAO observations play an important role in testing the CDDR Wu et al. (2015); Xu and Huang (2020); Wang et al. (2024); Xu et al. (2022). For example, Ref. Wu et al. (2015) tested the CDDR by comparing the SN data from Union 2.1 compilation with five BAO measurements, and concluded that the high precision of BAO measurements makes them an effective tool for validating the CDDR. Ref. Wang et al. (2024) tested the CDDR relation by utilizing SN data from the Pantheon sample and transverse BAO measurements from Sloan Digital Sky Survey (SDSS), demonstrating that transverse BAO measurements can serve as a powerful tool for validating the CDDR relation.
@@
Recently, the Dark Energy Spectroscopic Instrument (DESI) publicly released its second data release (DR2), which includes measurements of BAO from over 14 million extragalactic sources, spanning the redshift range $0.295\leq z\leq 2.330$. This release has marked the beginning of a new era in high-precision BAO measurements. The DESI DR2 BAO data, combined with SN data from PantheonPlus, Union3, and DESY5, along with CMB data from Planck and ACT, provide more stringent constraints on the dark energy equation of state, revealing a $2.8\sigma-4.2\sigma$ preference for dynamical dark energy Abdul Karim et al. (2025). DESI BAO data have attracted a significant number of researchers seeking to constrain various aspects of cosmological physics, such as dark energy Li et al. (2024); Giarè et al. (2024); Dinda and Maartens (2025); Escamilla et al. (2024); Sabogal et al. (2024); Li et al. (2025a); Li and Wang (2024); Wang and Piao (2024); Giarè et al. (2024); Huang et al. (2025b); Li et al. (2025b); Wu (2025); Li and Zhang (2025); Li and Wang (2025); Barua and Desai (2025); Yashiki (2025); Ling et al. (2025); Goswami and Das (2025); Yang et al. (2025a); Pang et al. (2025); You et al. (2025); Özülker et al. (2025); Cheng et al. (2025); Pan et al. (2025); Qiang et al. (2025); Liu et al. (2025a); Gialamas et al. (2025), non-cold dark matter Yang et al. (2025b); Chen and Loeb (2025); Wang (2025); Kumar et al. (2025); Abedin et al. (2025a); Khoury et al. (2025); Araya et al. (2025); Li et al. (2025c, d); Paliathanasis (2025), neutrino Du et al. (2025a); Jiang et al. (2025); Du et al. (2025b); Feng et al. (2025); Roy Choudhury (2025), and modified gravity Cai et al. (2025); Li et al. (2025e); Odintsov et al. (2025); Nojiri et al. (2025). Additionally, it is essential to revisit the CDDR in the context of the most recent high-precision BAO data from DESI, as emphasized by recent studies Zhang et al. (2025b); Wang et al. (2025); Santos et al. (2025); Afroz and Mukherjee (2025); Yang et al. (2025c); Keil et al. (2025); Teixeira et al. (2025). For example, Favale et al. (2024) investigated the consistency of the CDDR by combining SDSS and DESI mixed BAO data with SN data, reporting indications of a potential deviation at approximately the $2\sigma$ level. The authors attributed this deviation to possible tensions between the SN and BAO data, or potentially internal inconsistencies within the BAO data subsets.
@@
In this work, we test the CDDR using BAO data from DESI and SDSS, along with SN data from PantheonPlus and DESY5. We utilize the artificial neural network (ANN) approach to align the SN and BAO data at the same redshifts. Our aim is to utilize the pairwise independent combinations of the BAO and SN datasets to provide a comprehensive and robust analysis of the CDDR and investigate any potential violations of this fundamental relation.
@@
This work is organized as follows. In Sec. II, we describe the BAO and SN data used in our analysis and outline the methodology for testing the CDDR. In Sec. III, we report the constraint results and make some relevant discussions. The conclusion is given in Sec. IV.
@@
Any departure of $\eta(z)\neq 1$ at a given redshift signifies an inconsistency between the CDDR and the cosmological data. We parameterize $\eta(z)$ using three forms: (i) P1: $\eta(z)=1+\eta_{0}\,z$; (ii) P2: $\eta(z)=1+\eta_{0}\,z/(1+z)$; (iii) P3: $\eta(z)=1+\eta_{0}\,\ln(1+z)$.
@@
DR2 LRG1 0.51 13.59 $\pm$ 0.17 Abdul Karim et al. (2025)
DR2 LRG2 0.71 17.35 $\pm$ 0.18 Abdul Karim et al. (2025)
DR2 LRG3 + ELG1 0.93 21.58 $\pm$ 0.15 Abdul Karim et al. (2025)
DR2 ELG2 1.32 27.60 $\pm$ 0.32 Abdul Karim et al. (2025)
DR2 QSO 1.48 30.51 $\pm$ 0.76 Abdul Karim et al. (2025)
@@
BOSS LRG1 0.38 10.27 $\pm$ 0.15 Gil-Marín et al. (2017)
BOSS LRG2 0.51 13.38 $\pm$ 0.18 Gil-Marín et al. (2017)
eBOSS LRG 0.70 17.65 $\pm$ 0.30 Bautista et al. (2020)
eBOSS ELG 0.85 19.50 $\pm$ 1.00 Hou et al. (2020)
eBOSS QSO 1.48 30.21 $\pm$ 0.79 Hou et al. (2020)
@@
  • PantheonPlus: The PantheonPlus comprises 1550 spectroscopically confirmed supernovae from 18 different surveys, covering the redshift range $2.260.01<z<$ 2.26 Brout et al. (2022) [^2] Data available at https://github.com/PantheonPlusSH0ES/DataRelease.
  • DESY5: The DESY5 sample comprises 1635 photometrically classified supernovae from the released part of the full 5 yr data of the Dark Energy Survey collaboration (with the redshift range $1.30.1<z<$ 1.3), complemented by 194 low-redshift supernovae from the CfA3 Hicken et al. (2009), CfA4 Hicken et al. (2012), CSP Krisciunas et al. (2017), and Foundation Foley et al. (2018) samples (with redshifts in the range $0.10.025<z<$ 0.1), for a total of 1829 supernovae Abbott et al. (2024). [^3] Data available at https://github.com/des-science/DES-SN5YR.
@@
The distance measurements of SN crucially rely on the calibration of $M_{\rm B}$. For the PantheonPlus analysis, we firstly perform fits treating $M_{\rm B}$ as a free parameter. Additionally, to evaluate the impacts of different prior values of $M_{\rm B}$ on the CDDR test, we adopt two specific Gaussian priors: (i) $M_{\rm B}^{\rm D20}=-19.230\pm 0.040\,\mathrm{mag}$, from low‑$z$ SN ($0.150.023<z<$ 0.15) in $\Lambda$CDM via de‑marginalization of the SH0ES result Camarena and Marra (2020); Reid et al. (2019); (ii) $M_{\rm B}^{\rm B23}=-19.396\pm 0.016\,\mathrm{mag}$, from combined SN and BAO data Dinda and Banerjee (2023). Including the observational error in $M_{\rm B}$, the uncertainty in the distance modulus is $\sigma_{\mu}=\sqrt{\sigma_{M_{\rm B}}^{2}+\sigma_{m_{\rm B}}^{2}}\,$. Conversely, the DESY5 compilation provides $\mu$ and $\sigma_{\mu}$ directly, as $M_{\rm B}$ is already normalized within the calibration. Therefore, there is no prior on $M_{\rm B}$ is imposed for DESY5 analysis in this work.
For the measurement of $D_{\rm A}(z)$, we utilize the transverse comoving distance $D_{\mathrm{M}}(z)/r_{\mathrm{d}}$ derived from the BAO observations of the DESI DR2 Abdul Karim et al. (2025) and SDSS Gil-Marín et al. (2017); Hou et al. (2020); Bautista et al. (2020). The specific BAO measurements are listed in Table 1. The angular diameter distance $D_{\rm A}(z)$ is obtained through
@@
In this work, we adopt the sound horizon $r_{\mathrm{d}}=147.78~{}\mathrm{Mpc}$ corresponding to the Planck 2018 result Aghanim et al. (2020). It is also worth noting that, due to the lower redshift range of the DESY5 data, we do not include the DESI DR2 ELG2, QSO measurements, and the SDSS eBOSS QSO data point in the DESY5-related analysis.
@@
SDSS + PantheonPlus $0.040^{+0.130}_{-0.190}$ $-19.510^{+0.290}_{-0.230}$ $0.290^{+0.360}_{-0.810}$ $-19.630^{+0.580}_{-0.400}$ $0.100^{+0.220}_{-0.390}$ $-19.540^{+0.400}_{-0.300}$
DESI + PantheonPlus $0.060^{+0.120}_{-0.210}$ $-19.560^{+0.420}_{-0.290}$ $0.480^{+0.440}_{-0.990}$ $-19.810^{+0.790}_{-0.570}$ $0.170^{+0.220}_{-0.480}$ $-19.650^{+0.570}_{-0.390}$
SDSS + PantheonPlus ($M_{\rm B}^{\rm D20}$) $-0.120\pm 0.067$ $-19.230\pm 0.040$ $-0.240\pm 0.130$ $-19.230\pm 0.040$ $-0.178\pm 0.093$ $-19.230\pm 0.040$
DESI + PantheonPlus ($M_{\rm B}^{\rm D20}$) $-0.098\pm 0.052$ $-19.230\pm 0.040$ $-0.220\pm 0.113$ $-19.230\pm 0.040$ $-0.162\pm 0.068$ $-19.230\pm 0.040$
SDSS + PantheonPlus ($M_{\rm B}^{\rm B23}$) $-0.040\pm 0.071$ $-19.396\pm 0.016$ $-0.052\pm 0.131$ $-19.396\pm 0.016$ $-0.050\pm 0.102$ $-19.396\pm 0.016$
DESI + PantheonPlus ($M_{\rm B}^{\rm B23}$) $-0.035\pm 0.056$ $-19.396\pm 0.016$ $-0.070\pm 0.124$ $-19.396\pm 0.016$ $-0.043\pm 0.088$ $-19.396\pm 0.016$
SDSS + DESY5 $-0.035\pm 0.074$- $-0.080\pm 0.120$- $-0.059\pm 0.093$-
DESI + DESY5 $-0.052\pm 0.066$- $-0.090\pm 0.110$- $-0.068\pm 0.085$-

### v2

@@ Abstract
One of the most fundamental relationships in modern cosmology is the cosmic distance duality relation (CDDR), which describes the relationship between the angular diameter distance ($D_{\rm A}$) and the luminosity distance ($D_{\rm L}$), and is expressed as $\eta(z)=D_{\rm L}(z)(1+z)^{-2}/D_{\rm A}(z)=1$. In this work, we conduct a comprehensive test of the CDDR by combining baryon acoustic oscillation (BAO) data from the SDSS and DESI surveys with type Ia supernova (SN) data from PantheonPlus and DESY5. We utilize an artificial neural network approach to match the SN and BAO data at the same redshift. To explore potential violations of the CDDR, we consider three different parameterizations: (i) $\eta(z)=1+\eta_{0}z$; (ii) $\eta(z)=1+\eta_{0}z/(1+z)$; (iii) $\eta(z)=1+\eta_{0}\ln(1+z)$. Our results indicate that the calibration of the SN absolute magnitude $M_{\rm B}$ plays a crucial role in testing potential deviations from the CDDR, as there exists a significant negative correlation between $\eta_{0}$ and $M_{\rm B}$. For PantheonPlus analysis, when $M_{\rm B}$ is treated as a free parameter, no evidence of CDDR violation is found. In contrast, fixing $M_{\rm B}$ to the $M_{\rm B}^{\rm D20}$ prior with $19.230\pm 0.040$ mag leads to a deviation at approximately the $2\sigma$ level, while fixing $M_{\rm B}$ to the $M_{\rm B}^{\rm B23}$ prior with $19.396\pm 0.016$ mag remains in agreement with the CDDR. Furthermore, overall analyses based on the SDSS+DESY5 and DESI+DESY5 data consistently show no evidence of deviation from the CDDR.
@@ I Introduction
The cosmic distance duality relation (CDDR), also known as the Etherington relation Etherington1933, is a fundamental relationship in modern cosmology. This relationship is predicated on three fundamental assumptions: the description of spacetime by the metric theory of gravity, the propagation of photons along null geodesics, and the conservation of the photon number Bassett:2003vu. Specifically, for a given redshift, the relation can be expressed as $D_{\rm L}(z)=(1+z)^{2}D_{\rm A}(z)$, which establishes a connection between two critical measures of cosmic distance for an astronomical object: the luminosity distance ($D_{\rm L}$) and the angular diameter distance ($D_{\rm A}$). Therefore, testing the validity of the CDDR is essential for verifying the theories of gravity and the foundations of modern cosmology Avgoustidis2010; Holanda2011.
@@
In recent years, a considerable amount of research has employed various astronomical observations to test the CDDR Uzan:2004my; DeBernardis:2006ii; Li:2011exa; Nair:2011dp; Holanda:2010vb; Liang:2011gm; Meng:2011nt; Khedekar:2011gf; Goncalves:2011ha; Lima:2011ye; Cardone:2012vd; Yang:2013coa; Wu:2015ixa; Ma:2016bjt; Holanda:2016msr; More:2016fca; Rana:2017sfr; Li:2017zrx; Lin:2018qal; Qi:2019spg; Holanda:2019vmh; Zhou:2020moc; Qin:2021jqy; Bora:2021cjl; Mukherjee:2021kcu; Liu:2021fka; Xu:2022zlm; Tonghua:2023hdz; Yang:2024icv; Tang:2024zkc; Jesus:2024nrl; Qi:2024acx; Alfano:2025gie. Type Ia supernova (SN) have been widely used to determine the luminosity distance, whereas the angular diameter distance, is typically obtained through multiple observational techniques, including (but not limited to) the Sunyaev–Zeldovich effect and gas mass fraction measurements in galaxy clusters Holanda:2010vb; Li:2011exa, strong gravitational lensing systems Liao:2022gde; Ruan:2018dls; Qi:2024acx, the angular size of ultra-compact radio sources Li:2017zrx, and baryon acoustic oscillations (BAO) Wu:2015ixa. In particular, SN and BAO observations play an important role in testing the CDDR Wu:2015ixa; Xu:2020fxj; Wang:2024rxm; Xu:2022zlm. For example, Ref. Wu:2015ixa tested the CDDR by comparing the SN data from the SCP Union2.1 compilation with five BAO measurements, and concluded that the high precision of BAO measurements makes them an effective tool for validating the CDDR. Reference Wang:2024rxm tested the CDDR relation by utilizing SN data from the Pantheon sample and transverse BAO measurements from the Sloan Digital Sky Survey (SDSS), demonstrating that transverse BAO measurements can serve as a powerful tool for validating the CDDR relation.
@@
Recently, the Dark Energy Spectroscopic Instrument (DESI) publicly released its second data release (DR2), which includes measurements of BAO from over 14 million extragalactic sources, spanning the redshift range $0.295\leq z\leq 2.330$. This release has marked the beginning of a new era in high-precision BAO measurements. The DESI DR2 BAO data, combined with SN data from PantheonPlus, Union3, and DESY5, along with CMB data from Planck and ACT, provide more stringent constraints on the dark energy equation of state, revealing a $2.8\sigma-4.2\sigma$ preference for dynamical dark energy DESI:2025zgx. DESI BAO data have attracted a significant number of researchers seeking to constrain various aspects of cosmological physics, such as dark energy Li:2024qso; Giare:2024gpk; Dinda:2024ktd; Colgain:2024xqj; Escamilla:2024ahl; Colgain:2024ksa; Sabogal:2024yha; Colgain:2024mtg; Li:2024qus; Li:2024bwr; Wang:2024dka; Giare:2024smz; Huang:2025som; Li:2025owk; Wu:2025wyk; Li:2025ula; Li:2025ops; Barua:2025ypw; Yashiki:2025loj; Ling:2025lmw; Goswami:2025uih; Yang:2025boq; Colgain:2025nzf; Pang:2025lvh; You:2025uon; Ozulker:2025ehg; Cheng:2025lod; Pan:2025qwy; Qiang:2025cxp; Liu:2025myr; Gialamas:2025pwv; Colgain:2025fct; Wu:2025vfs, non-cold dark matter Yang:2025ume; Chen:2025wwn; Wang:2025zri; Kumar:2025etf; Abedin:2025dis; Khoury:2025txd; Araya:2025rqz; Li:2025eqh; Li:2025dwz; Paliathanasis:2025xxm, neutrinos Du:2024pai; Jiang:2024viw; Du:2025iow; Feng:2025mlo; RoyChoudhury:2025dhe; Du:2025xes; Zhou:2025nkb, and modified gravity Cai:2025mas; Li:2025cxn; Odintsov:2025jfq; Nojiri:2025low. Additionally, it is essential to revisit the CDDR in the context of the most recent high-precision BAO data from DESI, as emphasized by several studies based on the first data release of DESI Favale:2024sdq; Yang:2025qdg; Keil:2025ysb; Teixeira:2025czm and more recent analyses using DESI DR2 data Zhang:2025qbs; Wang:2025gus; Santos:2025gjf; Afroz:2025iwo. For example, Ref. Favale:2024sdq investigated the consistency of the CDDR by combining SDSS and DESI mixed BAO data with SN data, reporting indications of a potential deviation at approximately the $2\sigma$ level. The authors attributed this deviation to possible tensions between the SN and BAO data, or potentially internal inconsistencies within the BAO data subsets.
However, a major challenge in testing the CDDR using the latest DESI DR2 data is the difficulty of obtaining matched angular diameter distance and luminosity distance at identical redshifts, as most observational datasets provide these quantities at different redshift points. To address this issue, several methods for reconstructing the distance have been proposed, such as linear and polynomial fitting Holanda:2015zpz and Gaussian process Hees:2014lfa. Specifically, Ref. Wang:2019vxv proposed a non-parametric approach based on an artificial neural network (ANN) to reconstruct functions directly from observational data. In recent years, several studies have applied the ANN method to test the CDDR; see, e.g., Refs. Liu:2021fka; Xu:2022zlm; Tang:2022ykd; Qi:2024acx; Yang:2025qdg.
@@
In this work, our primary motivation is to test the validity of the CDDR by utilizing the latest DESI DR2 BAO and DESY5 SN data and by applying an ANN approach to align the BAO and SN data at the same redshifts. To obtain more comprehensive and robust results, we also incorporate the SDSS BAO and PantheonPlus SN data, and perform cross-checks using four pairwise independent combinations of the BAO and SN datasets.
@@
This work is organized as follows. In Sect. II, we describe the BAO and SN data used in our analysis and outline the methodology for testing the CDDR. In Sect. III, we report the constraint results and make some relevant discussions. The conclusion is given in Sect. IV.
@@ II Methodology and data
Any departure of $\eta(z)\neq 1$ at a given redshift signifies an inconsistency between the CDDR and the cosmological data. We parameterize $\eta(z)$ using three forms: (i) $P_{1}$: $\eta(z)=1+\eta_{0}\,z$; (ii) $P_{2}$: $\eta(z)=1+\eta_{0}\,z/(1+z)$; (iii) $P_{3}$: $\eta(z)=1+\eta_{0}\,\ln(1+z)$. The first function represents the first-order term of a Taylor series expansion, but it exhibits undesirable behavior at high redshift. In contrast, the second function performs well even at high redshift, evolving more gradually than the first. The third function is expressed in a more complex logarithmic form. The reason for considering these three typical parameterizations is that they are widely used in many studies (see, e.g., Refs. Liu:2021fka; Xu:2022zlm; Qi:2024acx; Yang:2024icv), facilitating a comparison with our results. Furthermore, it is important to highlight that the motivation for considering multiple parameterizations stems from previous studies, which show that in any expansion of the form $\eta(z)=a+bf(z)$, the associated uncertainties are highly sensitive to the choice of the function $f(z)$ (see, e.g., Ref. Colgain:2021pmf). This inherent arbitrariness in a single parameterization is best addressed by considering multiple parameterizations, thus enabling a more robust and unbiased analysis.
@@
DR2 LRG1 0.51 13.59 $\pm$ 0.17 DESI:2025zgx
DR2 LRG2 0.71 17.35 $\pm$ 0.18 DESI:2025zgx
DR2 LRG3 + ELG1 0.93 21.58 $\pm$ 0.15 DESI:2025zgx
DR2 ELG2 1.32 27.60 $\pm$ 0.32 DESI:2025zgx
DR2 QSO 1.48 30.51 $\pm$ 0.76 DESI:2025zgx
@@
BOSS LRG1 0.38 10.27 $\pm$ 0.15 Gil-Marin:2016wya
BOSS LRG2 0.51 13.38 $\pm$ 0.18 Gil-Marin:2016wya
eBOSS LRG 0.70 17.65 $\pm$ 0.30 eBOSS:2020lta
eBOSS ELG 0.85 19.50 $\pm$ 1.00 eBOSS:2020gbb
eBOSS QSO 1.48 30.21 $\pm$ 0.79 eBOSS:2020gbb
@@
  • PantheonPlus: The PantheonPlus comprises 1550 spectroscopically confirmed supernovae from 18 different surveys, covering the redshift range $0.01<z<2.26$ Brout:2022vxf. [^2] https://github.com/PantheonPlusSH0ES/DataRelease.
  • DESY5: The DESY5 sample comprises 1635 photometrically classified supernovae from the released part of the full 5-year data of the Dark Energy Survey collaboration (with a redshift range $0.1<z<1.3$), complemented by 194 low-redshift supernovae from the CfA3 Hicken:2009df, CfA4 Hicken:2012zr, CSP Krisciunas:2017yoe, and Foundation Foley:2017zdq samples (with redshifts in the range $0.025<z<0.1$), for a total of 1829 supernovae DES:2024jxu. [^3] https://github.com/des-science/DES-SN5YR.
@@
[…]
For the measurement of $D_{\rm A}(z)$, we utilize the transverse comoving distance $D_{\mathrm{M}}(z)/r_{\mathrm{d}}$ derived from the BAO observations of the DESI DR2 DESI:2025zgx and SDSS Gil-Marin:2016wya; eBOSS:2020gbb; eBOSS:2020lta. The specific BAO measurements are listed in Table 1. The angular diameter distance $D_{\rm A}(z)$ is obtained through
@@
In this work, we adopt the sound horizon $r_{\mathrm{d}}=147.78~\mathrm{Mpc}$ corresponding to the Planck 2018 result Planck:2018vyg. Note that we do not consider the uncertainty in $r_{\rm d}$, since it is much smaller than the uncertainties in $D_{\rm L}(z)$ and $D_{\rm A}(z)$ and is therefore negligible. Additionally, some DESI and SDSS BAO measurements are excluded from this analysis because their redshifts lie beyond the range probed by the SN samples. Specifically, when using the PantheonPlus SN sample, we omit BAO measurements from the Lyman-$\alpha$ forest in DESI DR2 and SDSS eBOSS. When using the DESY5 SN sample, we omit the DESI DR2 ELG2, QSO, and Lyman-$\alpha$ forest measurements, as well as the SDSS eBOSS QSO and Lyman-$\alpha$ forest data points.
@@
SDSS + PantheonPlus $0.040^{+0.130}_{-0.190}$ $19.510^{+0.290}_{-0.230}$ $0.290^{+0.360}_{-0.810}$ $19.630^{+0.580}_{-0.400}$ $0.100^{+0.220}_{-0.390}$ $19.540^{+0.400}_{-0.300}$
DESI + PantheonPlus $0.060^{+0.120}_{-0.210}$ $19.560^{+0.420}_{-0.290}$ $0.480^{+0.440}_{-0.990}$ $19.810^{+0.790}_{-0.570}$ $0.170^{+0.220}_{-0.480}$ $19.650^{+0.570}_{-0.390}$
SDSS + PantheonPlus $\times M_{\rm B}^{\rm D20}$ $0.120\pm 0.067$ $19.230\pm 0.040$ $0.240\pm 0.130$ $19.230\pm 0.040$ $0.178\pm 0.093$ $19.230\pm 0.040$
DESI + PantheonPlus $\times M_{\rm B}^{\rm D20}$ $0.098\pm 0.052$ $19.230\pm 0.040$ $0.220\pm 0.113$ $19.230\pm 0.040$ $0.162\pm 0.068$ $19.230\pm 0.040$
SDSS + PantheonPlus $\times M_{\rm B}^{\rm B23}$ $0.040\pm 0.071$ $19.396\pm 0.016$ $0.052\pm 0.131$ $19.396\pm 0.016$ $0.050\pm 0.102$ $19.396\pm 0.016$
DESI + PantheonPlus $\times M_{\rm B}^{\rm B23}$ $0.035\pm 0.056$ $19.396\pm 0.016$ $0.070\pm 0.124$ $19.396\pm 0.016$ $0.043\pm 0.088$ $19.396\pm 0.016$
SDSS + DESY5 $035\pm 0.074 - 0.0$$80\pm 0.120 - 0.0$$059\pm 0.093 -$
DESI + DESY5 $052\pm 0.066 - 0.0$$90\pm 0.110 - 0.0$$068\pm 0.085 -$
In order to match the BAO and SN data at a given redshift, we reconstruct $m_{\rm B}$ or $\mu$ of SN using a nonparametric ANN method implemented with the REFANN Python package Wang:2019vxv. This ANN method is fully data-driven, allowing us to reconstruct a function from any type of data without assuming a specific parameterization of the function. An ANN typically consists of three main layers: the input layer, the hidden layer, and the output layer. The input layer consists of the features of the dataset, with each neuron representing a feature or variable. Since our goal is to reconstruct the functional relationship from the real data, we use all available data as the training input, without retaining a separate testing set, which is consistent with many similar studies (see, e.g., Refs. Qi:2024acx; Yang:2024icv). The neurons in the hidden layer receive outputs from the previous layer and transmit them to the neurons in the next layer. The hidden layers in an ANN are responsible for learning and data processing, and can consist of one or more layers. In our analysis, a single hidden layer connects the supernova redshift with the corresponding $m_{\rm B}$ or $\mu$ values, with a total of 4096 neurons in the hidden layer. This choice of the number of neurons is consistent with our previous work, providing a reasonable balance between accuracy and computational efficiency Qi:2023oxv; Qi:2024acx. The output layer is primarily responsible for generating the final result. We use a nonlinear activation function (exponential linear unit), with the remaining hyperparameters set to their default values as specified in Ref. Wang:2019vxv, and the mean absolute error is used as the loss function. The optimization method employed is gradient descent, which iteratively updates the loss value in the opposite direction of the current gradient to reduce the loss. The training process is performed over 30000 iterations to ensure the convergence of the loss function.
The ANN method is trained on the PantheonPlus and DESY5 datasets to yield a data-driven reconstruction of $m_{\rm B}$ or $\mu$. It is worth noting that for the SN data, we used only the diagonal elements of the covariance matrix to facilitate reconstruction using ANN. From this reconstruction we extract $D_{\rm L}(z)$ at the redshifts of BAO measurements and compare it with the angular diameter distance $D_{\rm A}(z)$ inferred from the same BAO data to test the CDDR. The ANN reconstruction $m_{\rm B}$ for PantheonPlus and $\mu$ for DESY5 in this work are shown in Fig. 1. Furthermore, we construct the SN Hubble diagram and compute the distance duality ratio $\eta(z)$, together with their uncertainties at the corresponding BAO redshifts using Eqs. (1), (2), and (5). The detailed results and discussion are presented in Appendix A.
This nonparametric framework does not assume any predefined functional form, providing a flexible approach for probing potential deviations from the standard cosmological model Wang:2020dbt; Liu:2021fka; Qi:2023oxv; Liu:2024yib; Abedin:2025yru. The precision of the reconstructed $m_{\rm B}$ or $\mu$ and the width of their confidence intervals depend on hyperparameter choices, such as network architecture and training strategy, which influence the expressive power and generalization capability of ANN. Nonetheless, the ANN‑based reconstruction remains a powerful, model‑independent tool for evaluating the CDDR using BAO and SN data.
We perform Bayesian analysis using the publicly available package Cobaya [^4] https://github.com/CobayaSampler/cobaya. Torrado:2020dgo and assess the convergence of the Markov Chain Monte Carlo (MCMC) Lewis:2002ah; Lewis:2013hha chains using the Gelman-Rubin statistics quantity R1<0.01R-1<0.01 Gelman:1992zz. The MCMC chains are analyzed using GetDist [^5] https://github.com/cmbant/getdist. Lewis:2019xzd.
III Results and discussions